OpenStudy (mattyboyy):

Is there someone willing to help me with 15 questions to get me credit?

Mehek (mehek14):

15?!?!

OpenStudy (cookielate):

If you could also help me ??? if u can thats okay

OpenStudy (mattyboyy):

I'm kinda crappy at math but Ill try @cookielate and yeah theyre the last 15 questions @Mehek14

Mehek (mehek14):

ok I'll see how many I can help you with

OpenStudy (cookielate):

post the questions

OpenStudy (mattyboyy):

In the diagram, line AB and EF are horizontal lines and CB is a vertical line segment. If FB : FC = 4 : 3, what are the coordinates of point D? A) (-4,-2) B) (-5,-3) C) (-6,-4) D) (-7,-5) E) (-8,-6) @Mehek14 @cookielate

OpenStudy (cookielate):

(-6,-4)

OpenStudy (cookielate):

C

OpenStudy (mattyboyy):

Got it lol @cookielate Can I post the rest on here and just tag you? the question is closed thats why im asking

OpenStudy (cookielate):

okay you have 14 left yay

OpenStudy (mattyboyy):

13 lol I got the other one @cookielate

OpenStudy (cookielate):

oh yay

OpenStudy (mattyboyy):

Okay next question,

OpenStudy (cookielate):

okay

OpenStudy (mattyboyy):

Line AD, AF and DF are tangent to circle O through points C, B, and E, respectively. If measure of DAF=66 degrees and the measure of FDA=56 degrees what is the measure of EOB? A) 132 B) 102 C) 122 D) 112 E) 96

OpenStudy (mattyboyy):

@cookielate

OpenStudy (mattyboyy):

122?

OpenStudy (cookielate):

yes 122 because if you add 56 + 66 = 122 c:

OpenStudy (cookielate):

12 question

OpenStudy (mattyboyy):

The table contains the proof of the theorem of the relationship between slopes of parallel lines. What is the missing statement in step 2? A) BC/AB = DE/EF B) BC/EF = AB/DE C) BC = EF, AB = DE D) BC/AB = EF/DE E) BC x AB = -(EF x DE)

OpenStudy (mattyboyy):

Ill post the table

OpenStudy (cookielate):

so they both are congruent

OpenStudy (cookielate):

i think its a or d one of those

OpenStudy (mattyboyy):

These are the statements and (reasons) 1.Triangle ABC and DEF are similar (Given) 2. (MISSING) (Property of Similar Triangles) 3. This statement is intentionally left blank (Property of Proportion) 4. Slope of line AC = BC/AB 5. Slope of line DF = EF/DE (Both 4 & 5 are definition of slope) 6. Slope of line AC = slope of line DF (Substitution Property of Equality)

OpenStudy (mattyboyy):

@cookielate

OpenStudy (cookielate):

so what answer do u think it is

OpenStudy (mattyboyy):

A? I think

OpenStudy (mattyboyy):

No D I think lol @cookielate

OpenStudy (cookielate):

well its a because d letters are not suppose to be opposite

OpenStudy (mattyboyy):

So I was right the first time lol

OpenStudy (cookielate):

yes yay

OpenStudy (mattyboyy):

10 more

OpenStudy (mattyboyy):

A ray of light is reflected off a mirror such that the reflected ray is perpendicular to the original ray, as shown in the diagram. What is the equation of the reflected ray? A) y - x = 6 B) y + x = 6 C) -y + x = 6 D) y + x = 5 E) y - x = 5

OpenStudy (mattyboyy):

@cookielate

OpenStudy (cookielate):

its not c

OpenStudy (mattyboyy):

D? Maybe?

OpenStudy (cookielate):

also its not e

OpenStudy (cookielate):

why do you think its d

OpenStudy (mattyboyy):

http://openstudy.com/study#/updates/53502ee2e4b0285dde94a4b6 @cookielate

OpenStudy (mattyboyy):

@abbychandler

OpenStudy (anonymous):

what do you need

OpenStudy (mattyboyy):

Help with this question^

OpenStudy (mattyboyy):

@learningtime1234

OpenStudy (anonymous):

yeah. Youre right. its D

OpenStudy (cookielate):

xD

OpenStudy (mattyboyy):

Thanks, 31 of 40 left

OpenStudy (anonymous):

so 9 left? or do you have 31 left

OpenStudy (cookielate):

-faints

OpenStudy (anonymous):

ikr

OpenStudy (mattyboyy):

9 left lol

OpenStudy (anonymous):

oh okay. go for it

OpenStudy (cookielate):

i gtg do my 5 paragraphs is someone willing to help me

OpenStudy (anonymous):

what is it on

OpenStudy (cookielate):

:( :( :( :( :( :( :( :(

OpenStudy (anonymous):

i love english

OpenStudy (mattyboyy):

Yeah I can help with english i just at math lol

OpenStudy (anonymous):

do you have any more questions @mattyboyy ?

OpenStudy (cookielate):

thanks @mattyboyy

OpenStudy (mattyboyy):

In Triangle ABC, the angle bisectors meet at point D. Point E is on line AC, and DE is perpendicular to AC. Point F is the location where the perpendicular bisectors of the sides of the triangle meet. What is the radius of the largest circle that can fit inside Triangle ABC? A) AD B) BD C) DE D) DF E) EF

OpenStudy (mattyboyy):

Sometimes I have to type in the stuff by myself

OpenStudy (anonymous):

hold on let me draw this haha

OpenStudy (mattyboyy):

http://openstudy.com/study#/updates/535570d0e4b099894ae22679 The answers come up on the internet but I just want to be sure lol

OpenStudy (anonymous):

yea one sec

OpenStudy (mattyboyy):

@cookielate yay

OpenStudy (mattyboyy):

Its the same question @cookielate

OpenStudy (anonymous):

D

OpenStudy (mattyboyy):

Lol youre back

OpenStudy (cookielate):

what same question

OpenStudy (anonymous):

yeah i kmnow. sorry. i took an hour break xD

OpenStudy (mattyboyy):

Triangle PQR and triangle PSR are inscribed in circle A. What is the measure of angle QPR? A) 137 degrees B) 107 degrees C) 95 degrees D) 42 degrees E) 30 degrees

OpenStudy (anonymous):

E) 30 degrees i love geometry

OpenStudy (cookielate):

need two more paragraphs ughh -_-

OpenStudy (anonymous):

haha you got it!! =)

OpenStudy (mattyboyy):

What is the standard form of the equation of a circle that has its center at (-2, -3) and passes through the point (-2, 0)? A) (x + 2)^2 + (y +3)^2 = 9 B) (x - 2)^2 + (y - 3)^2 = 16 C) (x - 2)^2 + (y - 3)^2 = 4 D) (x + 2)^2 + (y + 3)^2 = 16 E) (x- 2)^2 + (y +3)^2 = 9

OpenStudy (anonymous):

id say B but im not sure on that one

OpenStudy (mattyboyy):

Lemme see if someone else knows for sure

OpenStudy (cookielate):

can someone help me on typing a paragraph please

OpenStudy (mattyboyy):

What is the equation of this circle in standard form? A) (x - 5.5)^2 + (y - 4)^2 = 7 B) (x + 5.5)^2 + (y + 4)^2 = 3.5 C) (x - 5.5)^2 + (y - 4)^2 = 3.5 D) (x + 5.5)^2 + (y + 4)^2 = 12.25 E) (x - 5.5)^2 + (y -4)^2 = 12.25

OpenStudy (cookielate):

if someone is nice enough or if u can

OpenStudy (anonymous):

yea. whats it on? and i would say C mattyboy

OpenStudy (mattyboyy):

alot of freaking typing -.-

OpenStudy (anonymous):

wait... actually i think D

OpenStudy (cookielate):

The equation of a circle in standard form is: (x−h)2+(y−k)2=r2 where the center is the point (h,k) and the radius is r. If we are given two end-points of a diameter (x1,y1) and (x2,y2), then we know the center of the circle must be the mid-point of the diameter, given by: (h,k)=(x1+x22,y1+y22) We also know the radius must be one-half the diameter, given by the distance formula as: r2=14((x2−x1)2+(y2−y1)2) Thus, the equation of the circle is: (x−x1+x22)2+(y−y1+y22)2=14((x2−x1)2+(y2−y1)2) Plugging in the given data: x1=2,y1=4,x2=9,y2=4 we find: (x−2+92)2+(y−4+42)2=14((9−2)2+(4−4)2) (x−112)2+(y−4)2=494=(72)2 This is equivalent to choice E.

OpenStudy (mattyboyy):

And yeah I mean I could post my question and type snippets of your essay, I kept the tab open lol

OpenStudy (mattyboyy):

Are you sure? @abbychandler lol

OpenStudy (cookielate):

really

OpenStudy (cookielate):

i think its e

OpenStudy (cookielate):

i even show you the example

OpenStudy (anonymous):

yea actually shes right. they should be negative. its E

OpenStudy (cookielate):

how many questions more

OpenStudy (mattyboyy):

5 questions left lol

OpenStudy (cookielate):

post all of them ill do three and @abbychandler could do two

OpenStudy (mattyboyy):

This diagram shows the aerial view of a playground. What is the area of the playground to the nearest square yard? Use the value pi = 3.14. A) 2,732 sq yrds B) 2,992 sq yrds C) 3,423 sq yrds D) 3,554 sq yrds E) 3,588 sq yrds